OpenStudy (anonymous):

A 13-foot ladder is leaning against a brick wall. The top of the ladder touches the wall 12 feet (ft) above the ground. The bottom of the ladder is 5 ft from the bottom of the wall. What is the sine of the angle formed by the ground and the base of the ladder?

OpenStudy (anonymous):

 And remember that \[\sin \theta = \frac{opposite}{hypotenuse}\]